consider two parents- one has Huntington’s disease (Hh), which is a dominant disorder, and the other does not. what is the probability that they will have a child with the disease?
a.0
b.1/4
c.1/2
d.3/4
e.1
Answer #1

Hh (affected) × hh (normal)

Gametes H h
h Hh (affected) hh (normal)

Proportion of affected child = 1/2

Percentage of affected child = 1/2 × 100 = 50%

Please rate high.